Circular No. 6255

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1995al IN NGC 3021 S. Pesci and P. Mazza, Milan, Italy, report their visual discovery of an

apparent supernova some 25′′ west-southwest of the center of NGC 3021. On Nov. 1.04 UT, they estimated m_{v} = 13.0–13.5, noting that no star appears at this position on the corresponding image in the Carnegie Atlas of Galaxies or on the Palomar Sky Survey. Pesci reports that nothing unusual was noted during an observation of this galaxy on Oct. 25.2, with an estimated stellar limiting visual mag of 16.0–16.2 outside of the galaxy (but perhaps closer to 15.0 in regions that are superimposed on the galaxy itself, as in this case). G. M. Hurst, Basingstoke, England, estimates mag 13.2 for SN 1995al from an unfiltered CCD image obtained by M. Armstrong (Rolvenden, U.K.) on Nov. 2.1.

C. Pollas, Observatoire de la Côte d’Azur (OCA), reports that a 10-min exposure with the OCA Schmidt telescope on Nov. 2.19 UT shows a star of mag V ∼ 14.3 that is not present on the original Palomar Sky Survey. SN 1995al is located at α = 9^{h}50^{m}55^{s}.97, δ = +33^{o}33′09′′.4 (equinox 2000.0), which is 15′′.0 west and 2′′.9 south of the center of NGC 3021. A nearby star (V ∼ 16.0) has position end figures 58^{s}.57, 27′′.40.

NOVA AQUILAE 1995 A. W. Neely, Silver City, NM, reports that the nova’s lightcurve in

V, R, and I continues a smooth, exponential decline, while showing a significant blueward shift. From Mar. 29 to Oct. 2, V − I dropped from +1.9 to +0.07. Magnitudes from Oct. 2: Johnson V = 13.12 ± 0.05; Cousins R = 12.92 ± 0.05; Cousins I = 13.05 ± 0.05.

COMET 122P/DE VICO Total visual magnitude estimates (cf. IAUC 6239): Oct. 1.13 UT, 5.4

(R. Haver, Frasso Sabino, Italy, 15×80 binoculars); 3.49, 5.2 (A. Hale, Cloudcroft, NM, naked eye); 5.48, 5.1 (R. Keen, Mt. Thorodin, CO, naked eye); 9.19, 5.3 (M. V. Zanotta, Laino, Italy, 20×80 binoculars); 15.11, 5.4 (L. Kiss, Szeged, Hungary, 20×60 binoculars); 19.23, 6.2 (A. Pereira, Cabo da Roca, Portugal, 9×34 binoculars); 22.17, 6.3 (A. Diepvens, Balen, Belgium, 8×50 binoculars); 28.74, 6.3 (H. Lüthen, Amrum, Germany, 10×50 binoculars); Nov. 2.09, 6.9 (C. E. Spratt, Victoria, BC, 0.10-m refractor).

1995 November 2 Daniel W. E. Green